OpenStudy (anonymous):

what is the inverse of 8 ^ -x ??

OpenStudy (anonymous):

also the how if possible

OpenStudy (anonymous):

1/8^x

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

umm its 8^x

OpenStudy (anonymous):

1/(8^-x)= 8^x

OpenStudy (anonymous):

how can (8^-x)=(8^x)

OpenStudy (anonymous):

its not. (8^-x)=1/(8^x) and then u want the inverse so u flip the fraction so it becomes 8^x